OpenStudy (anonymous):

what is the magnitude of the complex number z=-8+15i

OpenStudy (anonymous):

\[|a+bi|=\sqrt{a^2+b^2}\quad a,b\in \mathbb{R}\]

OpenStudy (anonymous):

huh? idk how to do this at all

OpenStudy (anonymous):

You take the numbers, square each one individually, add them together, and take the square root. Do each step one at a time.

OpenStudy (anonymous):

1. Take each number 2. Square each number 3. Add both of the squared numbers together. 4. Take the square root of the squares added together.

OpenStudy (anonymous):

@Limitless i have no idea what your saying to do, sorry !! this is like a pre unit lesson so i never learned this before

OpenStudy (zehanz):

If z=-8+15i, you should look upon it as the point (-8,15) in the complex plane. -8 and 15 are its coordinates. The distance of this point to the origin is z's magnitude. You can use Pythagorean Theorem to calculate this distance...

OpenStudy (anonymous):

Ok. First step: What are your numbers? Your numbers are \(8\) and \(15\). Second step: You square them. So: \(8^2\) and \(15^2\). Now what? Follow the other steps.

OpenStudy (anonymous):

@ZeHanz Introducing the complex plane might not be the best idea. It's directly identical, sure, but it assumes that the person has an understanding of ordered pairs, planes, magnitude, and Pythagorean theorem.

OpenStudy (zehanz):

@Limitless: I assumed knowlegde of complex numbers also means knowledge of the complex plane. In my country (Netherlands) only a few highschool students learn about complex numbers, but when they do, the complex plane is part of it...

OpenStudy (anonymous):

i do know about complez planes guys, thank you. the answer is 17

OpenStudy (zehanz):

YW!